# Differential Test Performance and Peer Effects

### v1

@@ 2.2 Peer Effects
We interpret $y_{crt}^{*}$ as the unobserved test taking ability in the absence of peer effects, but accounting for possible selection of student $r$ into classroom $c,$ taking test $t.$ The function $\psi_{crt}\left(.\right)$ accounts for the fact that test performance varies by student and thus depends on $r$ as well as additional student, teacher, classroom and school characteristics not captured by $y_{crt}^{*}$. Observed performance is determined by $y_{crt}=\psi_{crt}\left(y_{c1t}^{*},\ldots,y_{cn_{c}t}^{*}\right)$. The formulation of $\psi_{crt}\left(.\right)$ explicitly accounts for the possibility of peer effects by allowing for individual performance to depend on the characteristics of all peers. In the absence of peer effects the function $\psi_{crt}\left(.\right)$ simplifies to $y_{crt}=\psi_{crt}\left(y_{crt}^{*}\right).$ In other words, the performance of the $r$-th student in class $c$ only depends on own characteristics and class characteristics such as teachers and resources but not on the characteristics of other students in the class.
In empirical applications it may be difficult to work with a non-parametric framework accounting for peer effects. For this reason we focus on linear production functions with linear peer effects. Let $v^{c}$ and $w_{t}^{c},t=1,2$, be matrices with row dimension $n$ that capture observed test invariant and test varying class level characteristics not related to student characteristics, such as the class size, the gender, education and experience of the teacher for subject $t$, and let $X=\left(v^{p},w_{1}^{p},w_{2}^{p},v^{c},w_{1}^{c},w_{2}^{c}\right)$ denote the matrix of observations on all covariates. At the class level let $y_{ct}=\left(y_{c1t},\ldots,y_{cn_{c}t}\right)^{\prime}$, $y_{ct}^{*}=\left(y_{c1t}^{*},\ldots,y_{cn_{c}t}^{*}\right)^{\prime}$, $v_{c}^{p}=\left(v_{c1}^{p\prime},...,v_{cn_{c}}^{p\prime}\right)^{\prime}$, $w_{ct}^{p}=\left(w_{c1t}^{p\prime},...,w_{cn_{c}t}^{p\prime}\right)^{\prime}$ with similar definitions for $v_{c}^{c}$ and $w_{ct}^{c}$, and let $\psi_{ct}=\left(\psi_{c1t},\ldots,\psi_{cn_{c}t}\right)^{\prime}$. We impose the following restriction on the education production function.
@@
$y_{ct}=\psi_{ct}\left(y_{ct}^{*}\right)=\alpha_{c}f_{t}\mathbf{1}_{c}+X_{ct}^{c}\beta_{t}^{c}+\left(I_{c}+\rho M_{c}\right)\left(X_{ct}^{p}\beta_{t}^{p}+y_{ct}^{*}\right),$ (1)
where $\mathbf{1}_{c}=\left(1,....,1\right)^{\prime}$ is an $n_{c}\times 1$ vector and $M_{c}$ is some matrix of dimension $n_{c}\times n_{c}$ with zero diagonal elements and $\rho$ is a fixed parameter. The parameter $f_{t}$ captures test specific effects.
The part of the expression involving $M_{c}$ in (1) models peer effects. The vector $M_{c}y_{ct}^{*}$ contains weighted averages of the latent performance of classroom peers, and the parameter $\rho$ specifies the degree to which peer effects influence actual outcomes.  [^3] In the usual terminology of Cliff-Ord (1973,1981) models, $M_{c}y_{ct}^{*}$ is called a spatial lag and $\rho$ the corresponding parameter. The vector $M_{c}X_{ct}^{p}\beta_{t}^{p}$ represents contextual peer effects.
Linear specifications such as the one in Assumption 1 have been motivated as solutions to Nash games in the literature, see Calvó-Armengol, Patacchini, and Zenou (2009), Blume et al. (2015) or Pereda-Fernández (2017). Pereda-Fernández (2017) considers a full information game of student and teacher effort where student and teacher quality is known to the players and that leads to an equation like (1). Thus, Assumption 1 can be justified by assuming that student ability $y_{ct}^{*}$ is unobserved by the analyst but known to classmates. Calvó-Armengol, Patacchini, and Zenou (2009) consider a network game in effort under full information where agents have linear quadratic utility. Therefore the educational production function (1) can be understood as an optimal response function where $y_{ct}^{*}$ now has the interpretation of effort, unobserved by the analyst but known to peers. Finally, Blume et al. (2015) analyze a Bayes-Nash equilibrium of agents choosing effort and facing a linear quadratic utility function that depends on their own as well as their peer’s effort; cp. also Cohen-Cole, Liu, and Zenou (2018) and Drukker, Egger, and Prucha (2023). Agents maximize expected utility conditional on publicly observable characteristics as well as private information that is only known to the individual player but not the other agents. They argue, see Blume et al. (2015) p. 452, that their model is observationally equivalent to Calvó-Armengol, Patacchini, and Zenou (2009) as long as the row sums of $M_{c}$ are all equal, an assumption we impose below. In particular, as noted by Blume et al. (2015) p. 458, $y_{ct}^{*}$ then can be interpreted as ability, with agents expectations absorbed into a fixed effect. This can be matched to our setting by assuming that $y_{ct}^{*}$ is ability that is unobserved by the analyst.
In line with the spatial and social network literature building on the models introduced by Cliff-Ord (1973,1981), the matrix $M_{c}$ in Assumption 1 could be an arbitrary weight matrix describing peer interaction subject to certain measurability assumptions specified below. For exposition and concreteness we focus on the case where average characteristics of peers determine peer effects. Define
$M_{c}=\left(\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}-I_{c}\right)/\left(n_{c}-1\right)$
where $I_{c}$ is the $n_{c}\times n_{c}$ identity matrix. Then, $M_{c}$ is the operator that computes the leave out average of peer characteristics. Combined with the functional form in Assumption 1 this leads, using scalar notation, to an education production function $y_{crt}=\psi_{crt}\left(y_{ct}^{*}\right)$ for observed outcomes in terms of unobservables and observables as given below ($t=1,2$)
$y_{crt}=\alpha_{c}f_{t}+x_{ct}^{c}\beta_{t}^{c}+x_{crt}^{p}\beta_{t}^{p}+y_{crt}^{*}+\rho\left[(n_{c}-1)^{-1}\sum_{l=1,l\neq r}^{n_{c}}\left(x_{clt}^{p}\beta_{t}^{p}+y_{clt}^{*}\right)\right].$ (2)
2.3 Identification
In this paper we propose an identification strategy for the parameter $\rho$ that does not require randomization over $\mathscr{S}$ or variation in $S_{c}.$ Our argument proceeds in two steps. First, we propose a moment restriction on unobserved test taking ability $y_{it}^{*}$ and show that this restriction implies linear and quadratic moment restrictions for quantities we are able to observe in the data. We then show that $\rho$ can be identified by considering a quasi-differencing transformation of (2) and a GMM estimator that exploits the restrictions on the conditional mean and variance implied by our identifying assumption. The asymptotic theory for a general class of GMM estimators of this type was developed by Kuersteiner and Prucha (2020). Here, we rely on their theory to derive sharp identification results and inference procedures that are specific to this application.
We formalize the information structure of our model in the following definition. Recall that $\zeta^{\ast}$ stands for latent test performance, effort or ability prior to taking tests $t=1,2.$ We use the notation $-t$ in the definition below to denote the other test. We define the following:
@@ 2.3 Identification
Let $y_{-it}^{*}=\left(y_{1t}^{*},...,y_{i-1,t}^{*},y_{i+1,t}^{*},...,y_{n,t}^{*}\right)^{\prime}$ and $y_{t}^{*}=\left(y_{1t}^{*},...,y_{nt}^{*}\right)^{\prime}$. Define the sigma fields (information sets)
$\mathcal{F}_{n,i,t}$ $=\sigma\left(S,\alpha,X,z,\zeta^{\ast},y_{-t}^{\ast},y_{-i,t}^{*}\right)\text{\,for $t=1,2,$}$

### v2

@@ 2.2 Peer Effects
We interpret $y_{crt}^{*}$ as the unobserved test taking ability in the absence of peer effects, but accounting for possible selection of student $r$ into classroom $c,$ taking test $t.$ The function $\psi_{crt}\left(.\right)$ accounts for the fact that test performance varies by student and thus depends on $r$ as well as additional student, teacher, classroom and school characteristics not captured by $y_{crt}^{*}$. Observed performance is determined by $y_{crt}=\psi_{crt}\left(y_{c1t}^{*},\ldots,y_{cn_{c}t}^{*}\right)$. The formulation of $\psi_{crt}\left(.\right)$ explicitly accounts for the possibility of peer effects by allowing for individual performance to depend on the characteristics of all peers. In the absence of peer effects the function $\psi_{crt}\left(.\right)$ simplifies to $y_{crt}=\psi_{crt}\left(y_{crt}^{*}\right).$ In other words, the performance of the $r$-th student in class $c$ only depends on own characteristics and, for the generalized model presented below, on class characteristics such as teachers and resources but not on the characteristics of other students in the class.
At the class level let $y_{ct}=\left(y_{c1t},\ldots,y_{cn_{c}t}\right)^{\prime}$, $y_{ct}^{*}=\left(y_{c1t}^{*},\ldots,y_{cn_{c}t}^{*}\right)^{\prime}$, and let $\psi_{ct}=\left(\psi_{c1t},\ldots,\psi_{cn_{c}t}\right)^{\prime}$. We postulate the following education production function, which accounts for possible peer effects.
Assumption 1.
Assume that
$y_{ct}=\psi_{ct}\left(y_{ct}^{*}\right)=\alpha_{c}\mathbf{1}_{c}+\varphi_{c}+\left(I_{c}+\rho M_{c}\right)y_{ct}^{*},$ (1)
where $\mathbf{1}_{c}=\left(1,....,1\right)^{\prime}$ is an $n_{c}\times 1$ vector and $M_{c}$ is some matrix of dimension $n_{c}\times n_{c}$ with zero diagonal elements and $\rho$ is a fixed parameter. The vector $\varphi_{c}=\left(\varphi_{c1},...,\varphi_{cn_{c}}\right)^{\prime}$ contains measurement errors.
The part of the expression involving $M_{c}$ in (1) models peer effects. The vector $M_{c}y_{ct}^{*}$ contains weighted averages of the latent performance of classroom peers, and the parameter $\rho$ specifies the degree to which peer effects influence actual outcomes.  [^3] In the usual terminology of Cliff-Ord (1973,1981) models, $M_{c}y_{ct}^{*}$ is called a spatial lag and $\rho$ the corresponding parameter.
Peer effects specifications such as the one in Assumption 1 have been motivated as solutions to Nash games in the literature, see Calvó-Armengol, Patacchini, and Zenou (2009), Blume et al. (2015) or Pereda-Fernández (2017). Our model allows for measurement errors $\varphi_{c}$ affecting test scores. Consistent with our assumption that tests $t=1$ and $t=2$ measure essentially identical skills, measurement errors are assumed to be test invariant. We do not impose any restrictions on whether $\varphi_{c}$ is correlated with $y_{ct}^{*}$ or whether $\varphi_{c}$ is cross-sectionally correlated. The model in (1) is equivalent to $y_{ct}=\alpha_{c}\mathbf{1}_{c}+\left(I_{c}+\rho M_{c}\right)\left(y_{ct}^{*}+\varphi_{c}\right)$ where the measurement errors are in terms of ability $y_{ct}^{*}$ rather than outcomes $y_{ct}.$
In line with the spatial and social network literature building on the models introduced by Cliff-Ord (1973,1981), the matrix $M_{c}$ in Assumption 1 could be an arbitrary weight matrix describing peer interaction subject to certain measurability assumptions specified below. For exposition and concreteness we focus on the case where average characteristics of peers determine peer effects. Define $M_{c}=\left(\mathbf{1}_{c}\mathbf{1}_{c}^{\prime}-I_{c}\right)/\left(n_{c}-1\right)$ where $I_{c}$ is the $n_{c}\times n_{c}$ identity matrix. Then, $M_{c}$ is the operator that computes the leave out average of peer characteristics. Combined with the functional form in Assumption 1 this leads, using scalar notation, to an education production function $y_{crt}=\psi_{crt}\left(y_{ct}^{*}\right)$ for observed outcomes in terms of unobservables as given below ($t=1,2$)
$y_{crt}=\alpha_{c}+\varphi_{cr}+y_{crt}^{*}+\rho\left[(n_{c}-1)^{-1}\sum_{l=1,l\neq r}^{n_{c}}y_{clt}^{*}\right].$ (2)
2.3 Closely Related Skills
In this paper we propose an identification strategy for the parameter $\rho$ that does not require randomization over $\mathscr{S}$ or variation in $S_{c}.$ Our argument proceeds in two steps. First, we propose moment restrictions on unobserved test taking ability $y_{it}^{*}$ and show that these restrictions imply linear and quadratic moment restrictions for quantities we are able to observe in the data. We then show that $\rho$ can be identified by considering a quasi-differencing transformation of (2) and a GMM estimator that exploits the restrictions on the conditional mean and variance implied by our identifying assumption. The asymptotic theory for a general class of GMM estimators of this type was developed by Kuersteiner and Prucha (2020). Here, we rely on their theory to derive sharp identification results and inference procedures that are specific to this application.
We formalize the information structure of our model in the following definition. Recall that $\zeta^{*}=\left(\zeta_{1}^{*},\ldots,\zeta_{n}^{*}\right)^{\prime}=S\tilde{\zeta}^{\ast}$, which stands for unobserved individual characteristics, latent test performance, effort or ability prior to taking tests $t=1,2.$ We use the notation $-t$ in the definition below to denote the other test. We define the following:
Definition 2.1.
Let $\varphi=\left(\varphi_{1}^{\prime},...,\varphi_{C}^{\prime}\right)^{\prime}$, $y_{-it}^{*}=\left(y_{1t}^{*},...,y_{i-1,t}^{*},y_{i+1,t}^{*},...,y_{n,t}^{*}\right)^{\prime}$ and $y_{t}^{*}=\left(y_{1t}^{*},...,y_{nt}^{*}\right)^{\prime}$. Define the sigma fields (information sets) $\mathcal{F}_{n,i,t}=\sigma\left(S,\alpha,\varphi,\zeta^{\ast},y_{-t}^{\ast},y_{-i,t}^{*}\right)\text{\,for $t=1,2$}$. Let $\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,\varphi,\zeta^{\ast}\right)$ be the sigma field of all conditioning variables.
The information set $\mathcal{Z}_{n}^{*}$ consists of classroom allocations $S,$ as well as of unobserved variables $\zeta^{\ast}$, $\alpha$ and $\varphi.$ The following restriction on the conditional means of the latent outcomes is at the core of our proposed method of identifying peer effects. We note that the assumption maintains implicitly that the two test scores are measured on the same scale.
Assumption 2.
Assume that $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}$ where $\kappa_{i}$ is a random variable that is invariant to $t$ and is measurable with respect to $\mathcal{Z}_{n}^{*}$ In addition, assume that the random variable $y_{it}^{*}$ is either bounded, or there exists an $\eta>0$ and a random variable $y$ such that $\left|y_{it}^{*}\right|+\left|\kappa_{i}\right|\leq y$ with $E\left[\left|y\right|^{4+\eta}|\mathcal{Z}_{n}^{*}\right]\leq K_{y}<\infty$ for all $i$ and $t.$ In addition, $E\left[\left|\alpha_{c}\right|^{2+\eta}\right]\leq K_{\alpha}<\infty$ and $E\left[\left|\varphi_{cr}\right|^{2+\eta}\right]\leq K_{\varphi}<\infty$ for all $c$ and $r$.
The interpretation of the condition is that conditional on a hypothetical or actual baseline of test results $\zeta^{*}$, as well as information about group formation $S$, unobserved class level characteristics $\alpha$, as well as measurement errors $\varphi$ , additional test results do not change expected performance. A second implication of Assumption 2 is that expected performance does not depend on the specific test $t.$ In other words, $\kappa_{i}$ only varies by individual $i$ but not by test $t.$ It is in this sense that tests $t=1$ and $t=2$ measure essentially identical skills. Note that $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]=\kappa_{i}$ by iterated expectations and $\mathcal{Z}_{n}^{*}\subset\mathcal{F}_{n,i,t}.$ If $y_{it}^{*}$ is correlated with measurement errors $\varphi_{c}$ then this dependence is absorbed into $\kappa_{i}$. The restriction Assumption 2 imposes on the relationship between measurement error and $y_{it}^{*}$ is that any type of mean dependence is common to both tests $t$. A special case covered by Assumption 2 is that $y_{it}^{*}$ may be independent of measurement error.
The variable $\kappa_{i}$ is a function of the entire cross section of latent performance measures $\zeta^{*}$, of the allocation $S$ and unobserved characteristics that include classroom effects and measurement errors $\varphi$. It is generally cross-sectionally dependent in ways we do not restrict or specify and captures such common effects as the socioeconomic background of students and classmates, their family background including parental education and support, their interaction with peers, the exposure to their teachers and resources available in their classroom. It also depends on a student’s own ability. The restriction we impose is that these factors do not change between tests $t=1$ and $t=2$.
Baseline performance $\zeta^{*}$ does not necessarily have to be observed or realized and $\zeta^{*}$ can alternatively be interpreted as latent ability prior to taking the actual tests $t=1,2$. In this sense, $\kappa_{i}$ and $E\left[y_{it}^{*}|\mathcal{Z}_{n}^{*}\right]$ are also unobserved. A scenario where Assumption 2 is realistic arises when the same type of test is taken multiple times without additional training between iterations of the test, or when several tests are given that focus on related skills, as is the case in Project STAR data for grades K-3 and tests for reading, and word study skills. [^4] Another example of a test with multiple scores is the Comprehensive Testing Program (CTP) administered by the Educational Records Bureau and given to Grades 1-11. The CTP test consists of several main categories such as “Auditory Comprehension”, “Reading Comprehension” and “Mathematics” with each category consisting of additional subcategories that each receive separate scores. The assumption is less plausible if comparisons are attempted across different subjects in a high school or college setting, for tests given at different times during the school year, or measuring different skills in a professional or team setting.
The following theorem formally establishes a decomposition of $y_{it}^{*}$ into its conditional mean, $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$, that is dependent cross-sectionally as well as across tests, and into uncorrelated idiosyncratic noise $u_{it}$.
Theorem 2.1.
Suppose Assumption 2 holds. Define $u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=y_{it}^{*}-\kappa_{i},$ then by construction $E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0$ and $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{F}_{n,i,t}\right)=0$ for any $j\neq i$ or $t\neq s$. Furthermore, let $\mathcal{G}_{n,i,t}$ be any sigma field with $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$, then $E\left[u_{it}|\mathcal{G}_{n,i,t}\right]=0$ and $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{G}_{n,i,t}\right)=0$ for any $j\neq i$ or $t\neq s$, and $\operatorname{Cov}\left(u_{it},\kappa_{i}\right)=0$ etc. If $E\left[y_{it}^{*}-\kappa_{i}|\mathcal{G}_{n,i,t}\right]\neq 0$ for some $\mathcal{G}_{n,i,t}$, then Assumption 2 cannot hold.
Examples for the information sets $\mathcal{G}_{n,i,t}$ in the above theorem are $\mathcal{G}_{n,i,t}=\sigma\left(S,\alpha,\varphi,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*s}\right)$ where $y_{-i,t}^{*s}$ is a subvector of $y_{-i,t}^{*}$ or $\mathcal{G}_{n,i,t}=\mathcal{Z}_{n}^{*}.$ The above theorem makes clear that the main restriction we impose on $y_{it}^{*}$ is the assumption that latent test performance is not predictable conditional on observable characteristics and unobserved baseline ability. The properties $E\left[u_{it}|\mathcal{Z}_{n}^{*}\right]=0,$ $\operatorname{Cov}\left(u_{it},u_{js}|\mathcal{Z}_{n}^{*}\right)=0$ for any $j\neq i$ or $t\neq s$ and $\operatorname{Cov}\left(u_{it},\kappa_{i}\right)=0$ follow directly from Assumption 2 and are not additional assumptions imposed on the distribution of $y_{it}^{*}.$ In other words, Assumption 2 is necessary and sufficient for the decomposition of $y_{it}^{*}=\kappa_{i}+u_{it}$ with uncorrelated errors $u_{it}$. Note that $M_{c}$ is implicitly a function of class assignment $S$ and thus orthogonal to $u_{it}.$
We do not attach a specific economic or behavioral interpretation to the decomposition of $y_{it}^{*}$ into $\kappa_{i}$ and $u_{it}$. It is merely a statistical representation for the purpose of isolating variation in test scores that is correlated within groups but invariant across tests and variation that is idiosyncratic to the individual student and test. Let $\kappa_{c}=\left(\kappa_{c1},\ldots.\kappa_{cn_{c}}\right)^{\prime}$, $u_{ct}=(u_{c1t},\ldots,u_{cn_{c}t})^{\prime}$ and combine Assumptions 1 and 2 such that
$y_{ct}=\mu_{c}^{*}+\left(I_{c}+\rho M_{c}\right)u_{ct}$ (3)
with the unobserved class and individual effect defined as
$\mu_{c}^{\ast}=\alpha_{c}\mathbf{1}_{c}+\varphi_{c}+\left(I_{c}+\rho M_{c}\right)\kappa_{c}.$ (4)
By differencing (3), which eliminates $\mu_{c}^{\ast}$, we obtain
$y_{c1}-y_{c2}$ $=$ $\left(I_{c}+\rho M_{c}\right)(u_{c1}-u_{c2}).$ (5)
As seen from the above, differential measures of test scores purge performance measures from common unobserved influences captured by $\mu_{c}^{*}$ while still maintaining relevant information that can shed light on the existence of peer effects. Let $y_{t}=\left(y_{1t}^{\prime},\ldots,y_{Ct}^{\prime}\right)^{\prime}$ and $u_{t}=\left(u_{1t}^{\prime},\ldots,u_{Ct}^{\prime}\right)^{\prime}$ for $t=1,2$. In similar fashion define $M=\operatorname{diag}_{c=1}^{C}\left(M_{c}\right)$. Then we can write (5) for the entire sample more compactly as
$y_{1}=y_{2}+(I+\rho M)(u_{1}-u_{2}).$ (6)
Our framework is in contrast with much of the econometrics literature on the identification of peer effects where assumptions about the orthogonality of $u_{it}$ and $\kappa_{i}$ and covariates are usually imposed directly. These unobservable quantities often are given economic interpretations. Examples of papers with such assumptions include Lee (2007b) who assumes i.i.d. errors, Graham (2008) who assumes independent classroom and idiosyncratic errors conditional on class type, Arcidiacono et al. (2012) who assume uncorrelated idiosyncratic errors and fixed effects that are orthogonal to idiosyncratic errors, Rose (2017) who assumes uncorrelated idiosyncratic errors, Kuersteiner, Prucha, and Zeng (2023) who assume a random effects specification and Lewbel, Qu, and Tang (2023) who assume exogeneity of group formation. Often these assumptions are justified by random group selection. In our setting we neither assume that individuals are randomly assigned to groups, nor that they are randomly selected from a population or that they have otherwise observed or unobserved characteristics that are independent in the cross-section.
3 Estimation with Covariates
3.1 Model with Covariates
We extend the model in Section 2 by adding observable characteristics and by specifying the test average scores in a more flexible way. We collect test invariant and test varying student level characteristics for students in class $c$ in the matrices $v_{c}^{p}$ and $w_{ct}^{p},t=1,2$ with row dimension $n_{c}$. We also collect observed test invariant and test varying class level characteristics for class $c$ which are not related to student characteristics in the matrices $v_{c}^{c}$ and $w_{ct}^{c},t=1,2$, with row dimension $n_{c}$. Class level characteristics may include class size, the gender, education and experience of the teacher for subject $t$, etc. In the following let $v^{p}=\left(v_{1}^{p\prime},\ldots,v_{C}^{p\prime}\right)^{\prime}$, $w_{t}^{p}=\left(w_{1t}^{p\prime},\ldots,w_{Ct}^{p\prime}\right)^{\prime}$, $v^{c}=\left(v_{1}^{c\prime},\ldots,v_{C}^{c\prime}\right)^{\prime}$, $w_{t}^{c}=\left(w_{1t}^{c\prime},\ldots,w_{Ct}^{c\prime}\right)^{\prime}$, $t=1,2,$ and let $X=\left(v^{p},w_{1}^{p},w_{2}^{p},v^{c},w_{1}^{c},w_{2}^{c}\right)$ denote the matrix of observations on all covariates. Below is an extended version of Assumption 1.
@@ 3.1 Model with Covariates
$y_{ct}=\psi_{ct}\left(y_{ct}^{*}\right)=\alpha_{c}\mathbf{1}_{c}+\varphi_{c}+X_{ct}^{c}\beta_{t}^{c}+\left(I_{c}+\rho M_{c}\right)\left(X_{ct}^{p}\beta_{t}^{p}+y_{ct}^{*}\right),$ (7)
where $\mathbf{1}_{c}=\left(1,....,1\right)^{\prime}$ is an $n_{c}\times 1$ vector and $M_{c}$ is some matrix of dimension $n_{c}\times n_{c}$ with zero diagonal elements and $\rho$ is a fixed parameter.
The vector $M_{c}X_{ct}^{p}\beta_{t}^{p}$ represents contextual peer effects. We update the filtrations given in Definition 2.1 to include observable covariates.